Friday the 13th Owl shift summary (08/13/99)

Valeri Jejer (jejer@uvahea.phys.virginia.edu)
Fri, 13 Aug 1999 08:03:25 -0400

Shift Summary for owl shift, 13/Aug/99

On Shift: Jim G., Valeri.

Hours of beam available at all intensities: 3.5
Hours of beam available at nominal intensity: 3.0
Hours of E832 data (special runs not included): 2.8
Hours of special runs (laser scans, muon runs...): 0.3

Accounting of lost time:

switching from MUON to E832 mode (20 min of beam);
DA problems followed by a few drt_fresh_starts (12 min of beam).

========================================================================
Runs ended this shift:
Run 14213 - - LASER_SCAN
Run 14214 - - COSMIC
Run 14215 - - COSMIC
Run 14216 - 0.2 M evts - BEAM_MUON_4E12
Run 14216 - ??? - BEAM_832 (aborted and erased)
Run 13222 - in progress - BEAM_832
Plus a couple of cosmic runs after laser scan.

Tapes brought to FCC this shift: none
(including tapes for online split)

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Waiting for beam, strugling with DA, E832 datataking with a lot of
beam aborts.
========================================================================
Description of Shift:

No beam until 01:26, finished Laser Scan and two cosmic runs.

01:26 - 01:28 - Muon run.

Some DA problems, suspecting bad tape.

Smooth datataking after 4:00 with occasional beam aborts for a spill
or two.

run spill time comment
------------------------------------------------------------------------
14214 00:25 CsI 2847 has gain drift 10.47%.
14216 1 01:47 Ktevana Errors - missing FFFF word in Cr.53,
kfrend.
4 01:50 Lost beam due to TeV RF problems.
02:30 JAG cleaned up onsplit disk space.
02:51 Beam has returned for 2 spills.
02:53 Lost beam (a few aborts followed by TeV ramp
problems).
03:20 Run aborted - problem with ktev3_drv2.
03:30 Run init hung up on stage 2 -> drt_fresh_start.
03:40 Run init hung up again, this time on Stage 1 -
could not find EOT on ktev3_drv2. We pulled out
QKB089 and replaced it with new tape QKB090.
QKB089 contains Laser scan data.
03:42 Beam is back while we are in process of
drt_fresh_start. The muon run was overwritten since
we could not append to a tape with no label, the
data is still on onsplit disks for those who
interested.
03:52 Run 14220 died on startup w/ocprc error ->
drt_fresh_start.
03:56 Multiple beam aborts. SEM up to 8E12 for a couple of
spills. DCs survived.
04:00 Beam is back. We are still initializing the run.
14222 1 04:12 Finally managed to start run. Missing FFFF words
in crate 53. Kfrend between init and start didn't
help.
04:17 JAG does gas check - He tubes at low pressure
(400 psi), sends e-mail to DN.
04:37 DC online histograms show hit multiplicities above
limit: plane 4 - 4.9 and plane 9 - 5.04. (at 6:45
they are 4.82 and 4.98 respectively).
95 06:18 Missing FFFF word in crate 53 - kfrend.
========================================================================
Follow up Notes on Previous Shift

o do a muon run when beam comes back -- DONE.

o keep an eye on E/p for CsI channel 2847 -- DID.

========================================================================
Items to Follow up on Next Shift